# A Formal Analysis of Algorithms for Matroids and Greedoids

### v1

@@ Abstract
We present a formal analysis, in Isabelle/HOL [30], of optimisation algorithms for matroids, which are useful generalisations of combinatorial structures that occur in optimisation, and greedoids, which are a generalisation of matroids. Although some formalisation work has been done earlier on matroids, our work here presents the first formalisation of results on greedoids, and many results we formalise in relation to matroids are also formalised for the first time in this work. We formalise the analysis of a number of optimisation algorithms for matroids and greedoids. We also derive from those algorithms executable implementations of Kruskal’s algorithm for minimum spanning trees, an algorithm for maximum cardinality matching for bi-partite graphs, and Prim’s algorithm for computing minimum weight spanning trees.
@@
Matroids, Greedoids, Formal Verification, Combinatorial Optimisation, Isabelle/HOL
@@ 1 Introduction
Matroids are algebraic structures that generalise the concepts of graphs and matrices, where a matroid is a pair of sets $(E,\mathcal{F})$ satisfying a number of conditions, e.g. $E$ is a finite set and $\mathcal{F}\subseteq 2^{E}$. Although they have a generally rich mathematical structure which inherently justifies studying them, there are two main motivations to studying them in the context of combinatorial optimisation. First, optimisation problems defined on matroids generalise many standard combinatorial optimisation problems: e.g. the minimisation problem for matroids over modular objective functions generalises the travelling salesman problem, the shortest-path problem, minimum spanning trees, and Steiner trees, all of which are important standard optimisation problems. This means that verifying an algorithm for the minimisation problem, for instance, allows for deriving algorithms to solve many other problems (modulo some conditions that we will discuss later), and in many cases those derived algorithms are practical, albeit sometimes asymptotically slower than the fastest possible algorithms. The second reason to study matroids is that, since they are an abstraction of other concrete structures, like graphs, reasoning about (algorithms that process) them avoids a lot of the combinatorial reasoning and the case analyses encountered when reasoning about (algorithms processing) the more concrete structures.
@@
In this paper we formalise, in Isabelle/HOL [30], a number of results about matroids, and greedoids, which are a generalisation of matroids. We formally analyse greedy algorithms for the maximisation problems for both structures and prove guarantees on the optimality of the solutions those greedy algorithms find. We also formally analyse an algorithm for the matroid intersection problem, where one aims to find the largest member of $\mathcal{F}_{1}\cap\mathcal{F}_{2}$, for two matroids $(E,\mathcal{F}_{1})$ and $(E,\mathcal{F}_{2})$. We then instantiate the verified algorithms with concrete instances to obtain executable algorithms for standard graph problems: we obtain an $O(n^{2}\cdot\log^{2}n)$ implementation of Kruskal’s algorithm from the greedy algorithm for matroids, an $O(n\cdot m\cdot(\log n+\log m))$ algorithm for maximum cardinality bi-partite matching from the matroid intersection algorithm, and an $O(n\cdot m\log n)$ implementation of Prim’s algorithm for minimum weight spanning trees from the greedy algorithm for greedoids, where $n$ and $m$ are the number of vertices and edges of the input graph, respectively. [^1] These running times were not verified. In addition to verifying these algorithms, we formalise a number of results from the theory of matroids and greedoids, e.g. the fact that matroids can be fully characterised by the greedy algorithm, and the relationship between greedoids and antimatroids. With the exception of deriving Kruskal’s algorithm from the greedy algorithm for maximisation over matroids [14], all the above results were not formalised before in any theorem prover. In our formalisation, we follow Korte and Vygen’s textbook [21] and Schrijver’s textbook [33], with the former being our main reference.
Availability Our formalisation is available online at https://github.com/mabdula/Isabelle-Graph-Library.
@@ 2 Background
Matroids A set system $(E,\mathcal{F})$ consists of a carrier set $E$ and a collection of independent sets $\mathcal{F}\subseteq 2^{E}$. We only consider finite carriers. We call $(E,\mathcal{F})$ an independence system iff (M1) $\emptyset\in\mathcal{F}$ and (M2) $X\subseteq Y$ and $Y\in\mathcal{F}$ implies $X\in\mathcal{F}$. Members of $\mathcal{F}$ are called independent. Sets $F\subseteq E$ with $F\not\in\mathcal{F}$ are called dependent. A matroid is an independence system satisfying the augmentation property (M3): for $X,Y\in\mathcal{F}$ and $|X|>|Y|$, there is $x\in X\setminus Y$ with $Y\cup\{x\}\in\mathcal{F}$.
@@
Formalisation Keinholz already formalised independence systems, matroids, bases, circuits, lower ranks and basic properties of these in Isabelle/HOL [16] using locales. In Isabelle/HOL, locales [7] are named contexts allowing the fixing of constants and the assertion of assumptions on their properties, which will then be available within the locale for definitions and proofs. Keinholz gives locales indep_system and matroid shown in Listing LABEL:isabelle:indep_sys_matroid_defs, each of which fixes a carrier set carrier and an independence predicate indep specifying which subsets of carrier should be independent. We extend this formalisation with another equivalent definition of the upper rank. For the rank quotient, we explicitly define the argument of the minimum to be $1$ when $\rho(X)=r(X)$ to treat division by zero and empty carriers.
@@
Verification Methodology As mentioned earlier, a goal of this work is to formalise some of the theory of matroids and greedoids and then obtain executable algorithms to process them. We follow Nipkow’s [29] approach of using locales to model abstract data types (ADTs) and to implement data structures such as finite sets. The functions and assumptions of an ADT can be encapsulated into a locale. In that approach, one specifies invariants, which give the conditions for the data structure to be well-formed, and abstraction functions, which convert an instance of the data structure into a value of the abstract mathematical type which it represents (see Listing LABEL:isabelle:custom_set_specs, for instance). Implementing an algorithm defined using ADTs can be done by instantiating the ADTs with executable implementations. The approach thus allows correctness guarantees and a lot of the proofs to be done at an abstract mathematical level, while executable implementations are obtained.
@@ 2 $F\leftarrow\emptyset$; //initialise result
We define the state for the greedy algorithm as a record type, where carrier_list is a list of type ’a list consisting of the elements of the carrier set (corresponding to $E$) and result is the result set of type ’s which is constructed over the course of the algorithm (corresponding to $F$). With this state type, we can specify the algorithm as a function in Isabelle/HOL as shown in Listing LABEL:isabelle:indep_sys_specs.
@@ 3.2 Verification of the Best-In-Greedy Algorithm
Within the appropriate context in the locale Best_In_Greedy (which assumes the three axiom predicates described above and that c and order are valid), the lemma can be formulated as in (Listing LABEL:isabelle:best_in_greedy_lemma_1) in Isabelle/HOL. Here, c_set c S denotes the sum of the costs of the elements in a set S, where the cost function is c.
@@
Listing 6: The first correctness lemma for the Best-In-Greedy algorithm. t_set is the abstraction function for the set ADT.
@@
For the invariant preservation proofs for invar_4, defining some more auxiliary invariants and proving they are preserved across the algorithm is necessary. These are not part of the informal proof, since they are either fairly trivial (e.g. showing that result is always a subset of the current carrier prefix) or aspects specific to the formal proof (e.g. showing that result always satisfies the set ADT well-formedness invariant). Proving that invar_4 is preserved across the different execution paths of BestInGreedy then boils down to proving that if result is a basis of the current carrier prefix, the result set after one state update will still be a basis of the carrier prefix in the next step. We consider the two recursive cases of BestInGreedy, in which the current element is either added to result if it preserves independence, or left out otherwise. The two lemmas required for the invariant proofs in these two cases are formulated and proved in the context of the abstract indep_system locale. Through the theorems connecting the implementation of independence systems to the abstract theory of independence systems, we are able to use these theorems to complete the invariant preservation proofs of invar_4.
@@
In our formalisation, this lemma is stated with an additional existential quantifier for the order parameter. The proof in our formalisation proceeds similarly to the informal proof we follow, except that we explicitly construct the initial order list and use the stability of the sorting function to show that the sorting of the elements produces the ordering necessary for the proof. Additionally, proving that the greedy result is $B_{1}$ required defining some new loop invariants and showing that they are preserved across the algorithm.
@@
Instantiation for Spanning Forests and Oracles For an undirected graph $G$ with a set of edges $E$, a spanning tree $T$ is an acyclic subgraph connecting all vertices of $G$ to one another. A spanning forest is acyclic and connects all vertices that are connected via $E$.
The set of edges $E$ in undirected graphs forms a matroid under acyclicity, i.e. where the sets of edges forming acyclic components is the independence set. This is a so-called graphic matroid and it is a widely cited example of matroids  [21, 34]. Acyclicity of sets $X\subseteq E$ indeed satisfies the matroid axioms and a circuit would be a cycle in the ordinary sense. We formalise this in Isabelle/HOL, including the equivalence of being a basis w.r.t. acyclicity and being a spanning forest, making the greedy algorithm suitable to solve the minimum spanning forest problem correctly.
@@
We therefore have an extended locale that also specifies the behaviour of a weak oracle, which is that if (a) $X$ is independent and if (b) suitable data structure and auxiliary invariants are satisfied and (c) $x\not\in X$, then the oracle tells correctly whether $X\cup\{x\}$ is independent. Subsequently, we verify the greedy algorithm with a weak oracle and prove it equivalent to the first version when called on the empty set. As a result, correctness can be lifted. In the end, we obtain an executable algorithm for minimum spanning forests using a simple DFS as weak oracle. A simplified version of the correctness theorem for Kruskal’s Algorithm is stated in Listing LABEL:isabelle:kruskal_corr. The obtained implementation has a running time of $\mathcal{O}(n^{2}\cdot\log^{2}n)$.
@@ 3.3 Greedy Algorithm for Greedoids
Greedoids are a generalisation of matroids. Their definition (conditions (M1) and (M3)) is obtained by dropping the (M2) from the definition of matroids. Similar to matroids, we use a locale to define greedoids, fixing the carrier set, its family of independent subsets, and specifying the axioms of greedoids. Their intuition is derived from the edge set of undirected trees containing a fixed vertex $r$ of an undirected graph and its total set of edges.
Our formalisation of greedoids contains accessible set systems (a set system in which every independent set contains an element removing which the resulting set continues to be independent) and antimatroids (accessible set systems closed under union) as well as the relationships between greedoids, accessible set systems and antimatroids. This is used to prove the correctness of the greedy algorithm for greedoids. Note: we formalise set systems anew because the existing formalisation of independence systems [16] uses a finite carrier set along with the independence predicate. Their formalisation was designed to be extended to matroids using the augmentation property, which would not have worked for greedoids.
We now consider the optimisation of modular weight functions $c:2^{E}\rightarrow\mathbb{R}$ on greedoids. Modular weight functions satisfy $c(A\cup B)=c(a)+c(B)-c(A\cap B)$ for all $A,B\subseteq E$. The algorithm GreedoidGreedy keeps track of a current solution $X\in\mathcal{F}$, initially $\emptyset$. In every iteration, it searches $E$ in the order $e_{1},...,e_{n}$ for an $x$ s.t. $x\not\in X$ and $X\cup\{x\}\in\mathcal{F}$ (’candidate’). It takes the first candidate in $e_{1},...,e_{n}$ such that $c(\{x\})\geq c(\{y\})$ for all other candidates $y$ (’candidate search’). If there is no candidate $x$, the procedure terminates. Otherwise the first best candidate is inserted into $X$, followed by the next iteration. Similar to Theorem 3, one can characterise certain greedoids with that algorithm.
@@
First, we say that a greedoid $(E,\mathcal{F})$ has the strong exchange property (SEP) iff for all $A,\,B\,\in\,\mathcal{F}$, $B$ basis w.r.t $\mathcal{F}$, $A\subseteq B$ and $x\,\in E\setminus B$ with $A\cup\{x\}\in\mathcal{F}$, there is $y$ with $A\cup\{y\}\in\mathcal{F}$ and $(B-\{y\})\cup\{x\}\in\mathcal{F}$. GreedoidGreedy computes a maximum-weight basis in $\mathcal{F}$ for any order of iteration $e_{1},...,e_{n}$ and any modular cost function $c:2^{E}\rightarrow\mathbb{R}$ iff $(E,\mathcal{F})$ has the SEP.
For the second direction ($\Leftarrow$), the usual proof is by constructing a counterexample to show the contrapositive. This heavily depends on the order $e_{1},...,e_{n}$ for candidate search. In Isabelle/HOL (Listing LABEL:isabelle:greedoid), both costs and the order (formalised as a list) are fixed by a context in which the algorithm is modelled, making $e_{1},...,e_{n}$ an input to the algorithm, as well. For the sake of brevity we cannot give further details neither on the informal nor the formal proof and the reader may refer to the formalisation instead.
Executability In Isabelle/HOL, we give an executable function that is equivalent to the one in Listing LABEL:isabelle:greedoid. However, candidate search is not performed on all $e\in E$. For the current solution $X$, it only checks those $e$ where $e\in E\setminus X$ to find a best candidate.
Instantiation An arborescence $T$ around a vertex $r$ in a graph with edges $E$ is an acyclic, connected subgraph of $E$ that contains $r$. If $r$ is in the vertices on which edges in $E$ are incident, the set of arborescences around $r$ forms a greedoid. Bases are spanning trees of the connected component of $r$ in $E$. The arborescence greedoid has strong exchange, making the instantiation of GreedoidGreedy, namely, the Jarnik-Prim Algorithm [31] optimal. We obtain a simple $\mathcal{O}(n\cdot|E|\cdot\log n)$ implementation to compute maximum weight bases where $n$ is the number of vertices in the component around $r$.

### v2

@@ Abstract
We present a formal analysis, in Isabelle/HOL [30], of optimisation algorithms for matroids, which are useful generalisations of combinatorial structures that occur in optimisation, and greedoids, which are a generalisation of matroids. Although some formalisation work has been done earlier on matroids, our work here presents the first formalisation of results on greedoids, and many results we formalise in relation to matroids are also formalised for the first time in this work. We formalise the analysis of a number of optimisation algorithms for matroids and greedoids. We also derive from those algorithms executable implementations of Kruskal’s algorithm for computing optimal spanning trees, an algorithm for maximum cardinality matching for bi-partite graphs, and Prim’s algorithm for computing minimum weight spanning trees.
@@
Matroids, Greedoids, Combinatorial Optimisation, Graph Algorithms, Isabelle/HOL, Formal Verification
@@ 1 Introduction
Matroids are algebraic structures that generalise the concepts of graphs and matrices, where a matroid is a pair of sets $(E,\mathcal{F})$ satisfying a number of conditions, e.g. $E$ is a finite set and $\mathcal{F}\subseteq 2^{E}$. Although they have a generally rich mathematical structure which inherently justifies studying them, there are two main motivations to studying them in the context of combinatorial optimisation. First, optimisation problems defined on matroids generalise, albeit not always in a practically useful way, many standard combinatorial optimisation problems: e.g. the minimisation problem for matroids over modular objective functions generalises the travelling salesman problem, the shortest-path problem, minimum spanning trees, and Steiner trees, all of which are important standard optimisation problems. This means that verifying an algorithm for the minimisation problem, for instance, allows for deriving algorithms to solve many other problems (modulo some conditions that we will discuss later), and in many cases those derived algorithms are practical, albeit sometimes asymptotically slower than the fastest possible algorithms. The second reason to study matroids is that, since they are an abstraction of other concrete structures, like graphs, reasoning about (algorithms that process) them avoids a lot of the combinatorial reasoning and the case analyses encountered when reasoning about (algorithms processing) the more concrete structures.
@@
In this paper we formalise, in Isabelle/HOL [30], a number of results about matroids, and greedoids, which are a generalisation of matroids. We formally analyse greedy algorithms for the maximisation problems for both structures and prove guarantees on the optimality of the solutions those greedy algorithms find. We also formally analyse an algorithm for the matroid intersection problem, where one aims to find the largest member of $\mathcal{F}_{1}\cap\mathcal{F}_{2}$, for two matroids $(E,\mathcal{F}_{1})$ and $(E,\mathcal{F}_{2})$. We then instantiate the verified algorithms with concrete instances to obtain executable algorithms for standard graph problems: we obtain an $O(n^{2}\cdot\log^{2}n)$ implementation of Kruskal’s algorithm from the greedy algorithm for matroids, [^1] We verify a variant to compute maximum spanning trees. an $O(n\cdot m\cdot(\log n+\log m))$ algorithm for maximum cardinality bi-partite matching from the matroid intersection algorithm, and an $O(n\cdot m\cdot\log n)$ implementation of Prim’s algorithm for minimum weight spanning trees from the greedy algorithm for greedoids, where $n$ and $m$ are the number of vertices and edges of the input graph, respectively. [^2] We did not verify the running times of these algorithms in Isabelle/HOL. Comments on running time are an informal analysis only. In addition to verifying these algorithms, we formalise a number of results from the theory of matroids and greedoids, e.g. the fact that matroids can be fully characterised by the greedy algorithm, and the relationship between greedoids and antimatroids. With the exception of deriving Kruskal’s algorithm from the greedy algorithm for maximisation over matroids [15] and Edmonds’ Intersection Theorem (Theorem 5[28], all the above results were not formalised before in any theorem prover. In our formalisation, we follow Korte and Vygen’s textbook [21] and Schrijver’s textbook [33], with the former being our main reference.
@@ 2 Background
Matroids
A set system $(E,\mathcal{F})$ consists of a carrier set $E$ and a collection of independent sets $\mathcal{F}\subseteq 2^{E}$. We only consider finite carriers. We call $(E,\mathcal{F})$ an independence system iff (M1) $\emptyset\in\mathcal{F}$ and (M2) $X\subseteq Y$ and $Y\in\mathcal{F}$ implies $X\in\mathcal{F}$. Members of $\mathcal{F}$ are called independent. Sets $F\subseteq E$ with $F\not\in\mathcal{F}$ are called dependent. A matroid is an independence system satisfying the augmentation property (M3): for $X,Y\in\mathcal{F}$ and $|X|>|Y|$, there is $x\in X\setminus Y$ with $Y\cup\{x\}\in\mathcal{F}$.
@@
Formalisation
Keinholz already formalised independence systems, matroids, bases, circuits, lower ranks and basic properties of these in Isabelle/HOL [17] using locales. In Isabelle/HOL, locales [8] are named contexts allowing the fixing of constants and the assertion of assumptions on their properties, which will then be available within the locale for definitions and proofs. Keinholz gives locales indep_system and matroid shown in Listing LABEL:isabelle:indep_sys_matroid_defs, each of which fixes a carrier set carrier and an independence predicate indep specifying which subsets of carrier should be independent. We extend this formalisation with another equivalent definition of the upper rank. For the rank quotient, we explicitly define the argument of the minimum to be $1$ when $\rho(X)=r(X)$ to treat division by zero and empty carriers.
@@
Verification Methodology
As mentioned earlier, a goal of this work is to formalise some of the theory of matroids and greedoids and then obtain executable algorithms to process them. We follow Nipkow’s [29] approach of using locales to model abstract data types (ADTs) and to implement data structures such as finite sets. The functions and assumptions of an ADT can be encapsulated into a locale. In that approach, one specifies invariants, which give the conditions for the data structure to be well-formed, and abstraction functions, which convert an instance of the data structure into a value of the abstract mathematical type which it represents (see Listing LABEL:isabelle:custom_set_specs, for instance). Implementing an algorithm defined using ADTs can be done by instantiating the ADTs with executable implementations. The approach thus allows correctness guarantees and a lot of the proofs to be done at an abstract mathematical level, while executable implementations are obtained.
@@ 2 $F\leftarrow\emptyset$; //initialise result
We define the state for the greedy algorithm as a record type, where carrier_list is a list of type ’a list consisting of the elements of the carrier set (corresponding to $E$) and result is the result set of type ’s which is constructed over the course of the algorithm (corresponding to $F$). With this state type, we can specify the algorithm as a function in Isabelle/HOL as shown in Listing LABEL:isabelle:best_in_greedy_def.
@@ 3.2 Verification of the Best-In-Greedy Algorithm
Within the appropriate context in the locale Best_In_Greedy (which assumes the three axiom predicates described above and that c and order are valid), the lemma can be formulated as in Listing LABEL:isabelle:best_in_greedy_lemma_1 in Isabelle/HOL. Here, c_set c S denotes the sum of the costs of the elements in a set S, where the cost function is c.
@@
Listing 6: The first correctness lemma for the Best-In-Greedy algorithm. to_set is the abstraction function for the set ADT.
@@
For the invariant preservation proofs for invar_4, defining some more auxiliary invariants and proving they are preserved across the algorithm is necessary. These are not part of the informal proof, since they are either fairly trivial (e.g. showing that result is always a subset of the current carrier prefix) or aspects specific to the formal proof (e.g. showing that result always satisfies the set ADT well-formedness invariant). Proving that invar_4 is preserved across the different execution paths of BestInGreedy then boils down to proving that if result is a basis of the current carrier prefix, the result set after one state update will still be a basis of the carrier prefix in the next step. We consider the two recursive cases of BestInGreedy, in which the current element is either added to result if it preserves independence, or left out otherwise. The two lemmas required for the invariant proofs in these two cases are formulated and proven in the context of the abstract indep_system locale. Through the theorems connecting the implementation of independence systems to the abstract theory of independence systems, we are able to use these theorems to complete the invariant preservation proofs of invar_4.
@@
In our formalisation, this lemma is stated with an additional existential quantifier for the order parameter. The proof in our formalisation proceeds similarly to the informal proof we follow, except that we explicitly construct the initial order list and use the stability of the sorting function to show that the sorting of the elements produces the ordering necessary for the proof. Additionally, proving that the greedy result is $B_{1}$ requires defining some new loop invariants (for example that no element of $F-B_{1}$ is ever in the greedy result) and showing that they are preserved across the algorithm.
@@
Instantiation for Spanning Forests and Oracles
For an undirected graph $G$ with a set of edges $E$, a spanning tree $T$ is an acyclic subgraph connecting all vertices of $G$ to one another. A spanning forest is acyclic and connects all vertices that are connected via $E$.
The set of edges $E$ in undirected graphs forms a matroid under acyclicity, i.e. where those sets of edges that form acyclic subgraphs are independent. This is a so-called graphic matroid and it is a widely cited example of matroids  [21, 33]. Acyclicity of sets $X\subseteq E$ indeed satisfies the matroid axioms and a circuit would be a cycle in the ordinary sense. We formalise this in Isabelle/HOL, including the equivalence of being a basis w.r.t. acyclicity and being a spanning forest, making the greedy algorithm suitable to solve the maximum spanning forest problem correctly. Note that minimum spanning trees can be easily computed using this algorithm by adapting the input.
@@
We therefore have an extended locale that also specifies the behaviour of a weak oracle, which is that if (a) $X$ is independent and if (b) suitable data structure and auxiliary invariants are satisfied and (c) $x\not\in X$, then the oracle determines whether $X\cup\{x\}$ is independent. Subsequently, we verify the greedy algorithm with a weak oracle by proving it equivalent to the first version when called on the empty set. As a result, correctness can be lifted. In the end, we obtain an executable algorithm for maximum spanning forests using a simple DFS as weak oracle. A simplified version of the correctness theorem for Kruskal’s Algorithm is stated in Listing LABEL:isabelle:kruskal_corr. An informal analysis yields that this implementation has a running time of $\mathcal{O}(n^{2}\cdot\log^{2}n)$. $\mathcal{O}(|E|\cdot\iota)$ is the running time of the abstract algorithm where $\iota$ is the running time of the oracle.
@@ 3.3 Greedy Algorithm for Greedoids
Greedoids are a generalisation of matroids. Their definition (conditions (M1) and (M3)) is obtained by dropping the condition (M2) from the definition of matroids. Similar to matroids, we use a locale to define greedoids, fixing the carrier set, its family of independent subsets, and specifying the axioms of greedoids. Their intuition is derived from the edge set of undirected trees containing a fixed vertex $r$ of an undirected graph and its total set of edges. Due to space limits, we briefly survey results on greedoids in our library without expanding on proofs. Interested readers may refer to the formalisation instead for details.
We introduce notions of accessible set systems, which are set systems in which every independent set contains an element removing which the resulting set continues to be independent, and antimatroids, which are accessible set systems that are closed under union. We formalise two fundamental results [20, 21] that are necessary to verify the greedy algorithm (Theorem 4), which is the main result in our library on greedoids. Firstly, given an accessible set system $(E,\,\mathcal{F})$, for $X\,\in\,\mathcal{F}$, $|X|\,=\,k$, there exists an order $x_{1},x_{2},\,\ldots\,x_{k}$ for elements of $X$ such that $\forall\,i\,\leq\,k,\,\left\{x_{1},\ldots\,x_{i}\right\}\,\in\,\mathcal{F}$. Secondly, a greedoid $(E,\mathcal{F})$ is accessible, as stated in Section 14.1 of [21]. Thirdly, we prove that every antimatroid is a greedoid [21].
We also consider the optimisation of modular weight functions $c:2^{E}\rightarrow\mathbb{R}$ on greedoids. Modular weight functions satisfy $c(A\cup B)=c(A)+c(B)-c(A\cap B)$ for all $A,B\subseteq E$. The algorithm GreedoidGreedy (Listing LABEL:isabelle:greedoid) keeps track of a current solution $X\in\mathcal{F}$, initially $\emptyset$. In every iteration, it searches $E$ in the order $e_{1},...,e_{n}$ for a candidate $x$ s.t. $x\not\in X$ and $X\cup\{x\}\in\mathcal{F}$. It takes the first candidate in $e_{1},...,e_{n}$ such that $c(\{x\})\geq c(\{y\})$ for all other candidates $y$. If there is no candidate $x$, the procedure terminates. Otherwise the first best candidate is inserted into $X$, followed by the next iteration. Similar to Theorem 3, one can characterise certain greedoids with that algorithm.
@@
We say that a greedoid $(E,\mathcal{F})$ has the strong exchange property (SEP) iff for all $A,\,B\,\in\,\mathcal{F}$, $B$ basis w.r.t $\mathcal{F}$, $A\subseteq B$ and $x\,\in E\setminus B$ with $A\cup\{x\}\in\mathcal{F}$, there is $y$ with $A\cup\{y\}\in\mathcal{F}$ and $(B-\{y\})\cup\{x\}\in\mathcal{F}$. GreedoidGreedy computes a maximum-weight basis in $\mathcal{F}$ for any order of iteration $e_{1},...,e_{n}$ and any modular cost function $c:2^{E}\rightarrow\mathbb{R}$ iff $(E,\mathcal{F})$ has the SEP.
For the second direction ($\Leftarrow$), the usual proof is by constructing a counterexample to show the contrapositive. This heavily depends on the order $e_{1},...,e_{n}$ for candidate search, making use of the greedoids-accessibility relationship and order property of accessible set systems. In Isabelle/HOL (Listing LABEL:isabelle:greedoid), both costs and the order (formalised as a list) are fixed by a context in which the algorithm is modelled, making $e_{1},...,e_{n}$ an input to the algorithm, as well, similar to the order parameter for BestInGreedy.
We formalise set systems anew because the existing formalisation of independence systems [17] uses a finite carrier set along with the independence predicate and assumes (M2). Thus, a formalisation of greedoids, which has to drop (M2), cannot be obtained from that. Note that, although we prove that a matroid according to Keinholz’s formalisation is a greedoid, we formalise a distinct greedy-algorithm for matroids discussed in the last section rather than using GreedoidGreedy. This is because, even a matroid with the strong exchange property stated in Theorem 4, i.e. one for which GreedoidGreedy works, the worst-case running time of GreedoidGreedy is much worse than BestInGreedyGreedoidGreedy has a worst-case running time of $\mathcal{O}(|E|^{2}\cdot\iota)$, where $\iota$ is the running time of the independence oracle, while BestInGreedy has $\mathcal{O}(|E|\cdot\iota)$.
Executability
We give an executable function that is equivalent to the one in Listing LABEL:isabelle:greedoid. However, candidate search is not performed on all $e\in E$ since anything in $X$ will never be a next best candidate. For the current solution $X$, it only checks those $e$ where $e\in E\setminus X$.
Instantiation
An arborescence $T$ around a vertex $r$ in a graph with edges $E$ is an acyclic, connected subgraph of $E$ that contains $r$. If an edge in $E$ is incident on $r$, the set of arborescences around $r$ forms a greedoid. Bases are spanning trees of the connected component of $r$ in $E$. The arborescence greedoid has strong exchange, making the instantiation of GreedoidGreedy, namely, the Jarnik-Prim Algorithm [31] optimal. We obtain a $\mathcal{O}(n\cdot|E|\cdot\log n)$ implementation to compute maximum weight bases where $n$ is the number of vertices in the component around $r$. Prim’s and Kruskal’s algorithm demonstrate that bounds tighter than what can be deduced from the running time of the abstract algorithm, which is $\mathcal{O}(|E|^{2}\cdot\iota)$ and $\mathcal{O}(|E|\cdot\iota)$, respectively, is possible for concrete problems.